OpenStudy (anonymous):

Tell whether the pair of expressions are equivalent -11x and x(-11) Yes or No

OpenStudy (anonymous):

@agent0smith @satellite73

OpenStudy (anonymous):

that is the "commutative law" which says \[a\times b=b\times a\]

OpenStudy (anonymous):

@answers1233

OpenStudy (anonymous):

Oh gosh @satellite73 Im gonna need a bit more than that..lol

OpenStudy (adryin):

Yes

OpenStudy (anonymous):

it is always true, like \(2\times 8=8\times 2\) and \(-11\times x=x\times (-11)\)

OpenStudy (anonymous):

Oh okay so the answer is "Yes"

OpenStudy (anonymous):

in math they call that the "commutative property of multiplication"

OpenStudy (anonymous):

yes, the answer is "yes"

OpenStudy (anonymous):

Thank you both! @Adryin can I tag you in more?

OpenStudy (adryin):

Sure